# VisualCloze: A Universal Image Generation Framework via Visual In-Context Learning

### v2

@@ Abstract
[…]
@@ 1 Introduction
[…]
@@
Recent advancements in image generation, propelled by the progress of diffusion models [15, 88, 33], have led to a wide range of applications, including image editing [69], style transfer [64, 81], virtual try-on [11, 12], and personalized generation [38, 54], among others. However, these tasks typically require task-specific models, which limit efficiency and scalability for real-world applications. In recent years, there has been growing interest in universal generative models [39, 27, 44], aiming to handle diverse image generation tasks, even unseen ones, within a single unified framework. Despite significant progress, some critical issues remain to be addressed, such as (1) distinguishable and generalizable task instruction, (2) comprehensive task coverage during training, and (3) a unified model architecture.
@@
An ideal task instruction is crucial for guiding the model to process the desired task effectively. Existing methods primarily rely on language instructions [27, 44] or task-specific tokens [39] to distinguish the task to be performed. However, the complexity of visual tasks and the inherent gap between vision and language modalities make it hard for the model to understand language-only task descriptions, which leads to task confusion [39] and hinders generalization on unseen tasks [71, 35]. Moreover, pre-learned task-specific tokens constrain the model only to handle seen tasks. In contrast, large language models (LLMs) have successfully achieved unified multi-task modeling, partially due to the rise of in-context learning [5], which allows models to adapt various tasks using only a few demonstrations. We aim to replicate the concept of in-context learning in the pure visual modality, where the model learns the desired task directly from a few visual examples as task demonstrations, as shown in Fig. 1 (Left Top). In this setting, in-context learning shows strong potential for universal image generation. We summarize four key findings: (1) it supports various in-domain tasks with reduced task ambiguity (Fig. 1); (2) it generalizes to unseen tasks (Fig. 2, Fig. 8); (3) as an unseen strategy for task unification, it can integrate multiple sub-tasks into a single step and generate intermediate results (Fig. 3); (4) it enables reverse generation, i.e., inferring a set of conditions from a given target (Fig. 9). While prior works [71, 4, 82, 3, 66, 43, 1] have also explored in-context learning in vision, they are largely constrained to specific domains (such as dense prediction or style transfer [67, 87]), or simplified generation settings involving only one condition and one target image [43, 60].
@@
From the perspective of task distribution, visual tasks are inherently sparse compared to those in natural language processing because task-specific datasets [71, 85] for different tasks have minimal overlap [79, 32, 19]. Such sparse task learning isolates the knowledge of each task and limits the model from learning shared features across tasks. Moreover, the weak correlations between tasks hinder knowledge transfer and adaptability to new tasks. However, existing works in multi-task learning [53, 31, 16, 10] have verified the benefits of overlapping knowledge across related tasks. To alleviate the sparsity of visual tasks, we introduce a graph-structured dataset, Graph200K, where each image is associated with annotations spanning five meta-tasks, i.e., conditional generation [80], IP preservation [76], style transfer [81], image editing [69], and restoration [77]. By combining different conditions, we train the model with a variety of tasks that overlap with each other. Given this highly overlapping and compact task space, our dataset significantly increases task density, allowing the model to learn shared and transferable knowledge more effectively.
@@
For the architecture design, it is essential to 1) accommodate flexible task formats [27, 71, 35], ensuring seamless in-context learning, and 2) remain compatible with state-of-the-art models [88, 33] to fully leverage their strong generative priors. In this work, we find that the state-of-the-art image infilling model [33] has a consistent objective with our in-context learning based universal generative formulation. Specifically, we concatenate all input and output images together, where the objective of a task is to fill the output area. This alignment enables us to build our model upon advanced general-purpose infilling models without additional modifications, achieving powerful universal generation capabilities with minimal data and training costs.
@@
In this work, we propose a universal image generation framework, VisualCloze, which fine-tunes FLUX.1-Fill-dev [33] with interrelated tasks sampled from Graph200K to learn transferable knowledge and support visual in-context learning. As the number of in-context examples increases, we observe enhanced performances and reduced task confusion, enabling the model to support a broad spectrum of in-domain tasks, including conditional generation, image restoration, editing, style transfer, IP-preservation, and their combinations. On unseen tasks, the model also shows a certain degree of generalization ability, as shown in Fig. 2. In summary, our main contributions are as follows:
@@ 2.1 Image Generation
Recent advances in text-to-image generation have achieved remarkable performance, largely driven by the development of autoregressive models [58, 78, 41] and diffusion models [24, 18, 48, 13, 51, 15, 2, 40, 42]. Among these, rectified flow transformers [33, 15, 17, 88] have shown great training efficiency and overall performance. Building on these foundational models, diverse applications have emerged, such as conditional generation [80], style transfer [64], and personalized generation [38]. More recently, universal models that address various tasks [83, 44, 35] have been explored. For example, unified models like OmniGen [71] leverage large vision language models to consolidate multiple tasks into a single framework. Similarly, UniReal [9] unifies image generation tasks as discontinuous video generation. However, they still face issues such as over-reliance on language instructions, isolation and sparsity of visual tasks, and architecture design accommodating flexible task formats. To address these issues, we propose a universal image generation framework that unifies generation tasks as image infilling. Through visual in-context learning and our Graph200K dataset that constructs a denser task space to learn transferable knowledge, our method alleviates ambiguity to support a diverse set of in-domain tasks and generalizes to tasks unseen during training.
@@ 2.2 Visual In-context Learning
Along with the emergence of large language models, such as GPT-3 [5], in-context learning [14] has been an effective approach to allow the language model to understand and perform complex tasks given a few demonstrations. Early works [22, 21] in vision modality propose image analogies to create an image filter from examples automatically. In recent years, leveraging inpainting model [4, 82, 3], masked image modeling [67, 66, 43], or vision-language model [1, 86], visual in-context learning is proposed to handle more tasks. However, they mainly focus on dense prediction [87, 55, 59] or visual understanding [63]. OmniGen [71] also leverages in-context learning to generalize to unseen domains, e.g., segmenting unseen concepts when the model has learned the segmentation task during training. However, it mainly focuses on simple tasks of dense prediction, and the gap between the unseen and training domains is still limited. Some recent works [43, 60, 68, 34] extend visual in-context learning to image generation, but they are still limited by simple tasks such as conditional generation and dense prediction. Moreover, the sparsity of visual tasks makes it difficult for models to learn transferable and overlapping knowledge across tasks, limiting the generation ability of in-context learning. In contrast, we introduce a graph-structured dataset that supports interrelated tasks and thus constructs a more dense task space, promoting the model to learn shared and transferable knowledge and enhance its adaptability.
@@ 3 Dataset
Recent works [44, 26, 71] have made great progress in unified image generation. However, their generalization to unseen tasks remains highly limited. We partially attribute this issue to the sparsity and isolation of visual tasks, hindering the model from learning shared features across tasks and handling unseen ones. Moreover, weak correlations between tasks further hinder knowledge transfer, restricting the adaptability of models. Therefore, increasing task density or strengthening task inter-relations helps improve the generalization ability of models via a compact task distribution. In this paper, we take the Subject200K [61] dataset as a starting point and construct our Graph200K dataset by augmenting each image with 49 types of annotations spanning five meta-tasks. This enriched annotation space enables flexible construction of a wide range of related tasks by sampling and combining arbitrary subsets of annotations across different meta-tasks, as illustrated in Fig. 4.
@@ 3.2 Dataset Construction
For convenience, we inherit subject-driven data from the Subjects200K [61]. Additionally, 32 different degradations are applied online to the images to acquire restoration data. We summarize the data construction methods in this section for the remaining three tasks.
Conditional generation. Each image is paired with 12 distinct conditions generated by specialized models, including canny edges [6], HED edges [72], Hough lines [20], semantic segmentation maps [37], depth maps [74], shape normal maps [73], and human keypoints [7], following ControlNet [80]. This work extends the conditions by incorporating SAM2 [50] masks, foreground segmentation, and open-world boxes and masks. The foreground segmentation, derived from the RMBG [84], supports diverse tasks such as inpainting and foreground extraction. Open-world bounding boxes are generated through the grounding caption capability of Qwen2-VL [65], which are processed using SAM2 [50] to produce corresponding masks.
Style transfer. We transfer the style of images according to reference in both semantic-variant and semantic-invariant settings. Specifically, the semantic-invariant transfer adopts InstantStyle [64] to preserve the semantic content, while the semantic-variant transfer relies on FLUX.1-Redux-dev [33], using the style embeddings and depth as conditions. For each image, we randomly generate five stylized versions. Mixing the two tasks pushes the model to follow the in-context examples better to avoid ambiguity.
Image editing. We design two types of editing tasks, including background-variant and background-invariant editing. The background-invariant editing begins with localizing the subjects. Then, we leverage a large vision-language model, Qwen2-VL [65], to modify the image caption with a new object that replaces the original subject. The image, with the subject masked, is subsequently processed by the FLUX.1-Fill-dev [33] inpainting model to integrate the alternative object into the masked region. The above operation is repeated five times to enrich the dataset. For background-variant editing, the difference lies in the last step, which utilizes FLUX.1-Redux-dev [33] with depth as the condition and the modified caption as the text prompt.
@@ 3.3 Other Data
To further expand the range of tasks and enhance the generalization ability of models, we incorporate several open-source datasets during training, including VITON-HD [11] for virtual try-on and PhotoDoodle [28] for artistic image editing. For image editing tasks, we also extend the dataset with OmniEdit [69]. Specifically, two sub-tasks, i.e., object addition and removal, are used for training. The other editing tasks, such as attribute modification and environment change, are treated as unseen tasks to assess the generalization ability of the trained model. Furthermore, we leverage a portion of high-quality internal data, covering tasks of the drawing process [62] and multi-view generation [29].
@@ 4.1 Visual In-context Learning
Language instructions are usually used to specify the generation definition to handle multiple visual generation tasks with a single generative model. However, due to the gap between vision and language, the text comprehension ability of image generation models remains limited. This issue leads to task confusion [39] in existing universal generative models and weak generalization to unseen tasks. Inspired by the success of few-shot learning on large language models [5], we recognize that visual context may serve as a more friendly task instruction for visual generative models, given their superior visual understanding capabilities.
Therefore, in this paper, we re-propose visual in-context learning to build a universal and generalizable image generation system. For the sake of description, here we assume the image input-output of arbitrary conditional generation task as a query consisting of $L-1$ condition images and a blank target $\varnothing$ to be completed by the model, i.e., ${X}={\rm concat}(\{x_{1},\dots,x_{L-1},\varnothing\})$. In Sec. 5.1, we demonstrate that our method can be extended to more general scenarios, where it can generate images at arbitrary positions and in any quantity rather than just the single image at the end of the query. During training, we randomly provide up to $C$ in-context examples, each containing $L$ images as the query. This strategy ensures the generalization ability of models across different numbers of in-context examples. In our experiments, we show that providing in-context examples as task demonstrations not only helps alleviate task confusion and boost model performance across in-domain tasks [39], but also enhances the generalization ability on unseen tasks.
@@ 4.2 Unified Multi-task Framework
Unlike previous visual in-context learning methods that primarily focus on scenarios with a single image condition and a single context [43, 60], in this work, we aim to construct a unified framework capable of handling varying numbers of conditions and contexts, allowing for flexible adaptation to diverse tasks. For ease of description, we first assume all images processed by the model share the same size, $W\times H$, and we extend to the scenario with mismatched aspect ratios at the end of this section. In this way, given $C$ in-context examples and the query, each containing $L$ images, all images can be concatenated into a complete grid-layout image with a size of $(L\times W,(C+1)\times H)$. Then, the model can complete a task by infilling the target grids based on the surrounding context, akin to solving visual cloze puzzles. Therefore, we build our unified framework, VisualCloze, based on the general image infilling architecture capable of handling multiple resolutions.
@@
where $M(i,j)=1$ indicates that the pixel will be masked and generated by the infilling model. Equ. (2) masks the region in the last row and column, i.e., the target image. During training, we also randomly mask one of the first $L-1$ grids with a probability of 0.5, promoting reverse generation shown in Sec. 5.1. For the inference stage, we can crop $\hat{X}$ to obtain the target image easily.
Aligned optimization objective. A key benefit of this design is that our VisualCloze formulation shares a highly consistent objective with general image infilling models without architectural modifications or explicit input conditions. This consistency allows us to directly fine-tune advanced image infilling models using the newly constructed dataset while maximizing the utilization of the prior knowledge of foundation models. In contrast, existing task-specific models often require introducing additional learnable modules [38, 69] or adapting to extra condition inputs [61], which may compromise the native capabilities of the model.
@@
Positional embedding. In the preceding section, all images are concatenated into a grid-layout image and we can apply positional embedding (i.e., RoPE [57]) on this large image. However, a potential limitation lies in composing a grid image from in-context examples with varying aspect ratios. To overcome this issue, we leverage the 3D-RoPE in Flux.1-Fill-dev to concatenate the query and in-context examples along the temporal dimension, as shown in Fig. 5, effectively overcoming this issue without introducing any noticeable performance degradation.

### v3

@@ Abstract
Figure 2: Unseen Tasks U: Generalizing to tasks unseen during training via in-context learning. More in-context examples lead to more accurate results.
@@ 1 Introduction
Figure 3: Unseen Tasks U: Leveraging in-context learning to unify multiple seen tasks into a single-step unseen task. Left: Unifying the [Depth to Image] and [Relighting] task into a single [Depth to Images with Various Lighting] task. Right: Unifying multiple dense prediction tasks into a joint prediction task. Results without visual context can be found in the appendix.
@@
Recent advancements in image generation, propelled by the progress of diffusion models [esser2024scalingrectifiedflowtransformers, zhuo2024luminanext, flux2024], have led to a wide range of applications, including image editing [wei2024omniedit], style transfer [wang2024instantstyle, Zhang_2023_inst], virtual try-on [choi2021viton, chong2025catvton], and personalized generation [li2023photomaker, ruiz2023dreambooth], among others. However, these tasks typically require task-specific models, which limit efficiency and scalability for real-world applications. In recent years, there has been growing interest in universal generative models [lin2024pixwizard, lhhuang2024iclora, mao2025ace++], aiming to handle diverse image generation tasks, even unseen ones, within a single unified framework. Despite significant progress, some critical issues remain to be addressed, such as (1) distinguishable and generalizable task instruction, (2) comprehensive task coverage during training, and (3) a unified model architecture.
@@
An ideal task instruction is crucial for guiding the model to process the desired task effectively. Existing methods primarily rely on language instructions [lhhuang2024iclora, mao2025ace++] or task-specific tokens [lin2024pixwizard] to distinguish the task to be performed. However, the complexity of visual tasks and the inherent gap between vision and language modalities make it hard for the model to understand language-only task descriptions, which leads to task confusion [lin2024pixwizard] and hinders generalization on unseen tasks [xiao2024omnigen, le2024diffusiongenerate]. Moreover, pre-learned task-specific tokens constrain the model only to handle seen tasks. In contrast, large language models (LLMs) have successfully achieved unified multi-task modeling, partially due to the rise of in-context learning [brown2020language], which allows models to adapt various tasks using only a few demonstrations. We aim to replicate the concept of in-context learning in the pure visual modality, where the model learns the desired task directly from a few visual examples as task demonstrations, as shown in Fig. 1 (Left Top). In this setting, in-context learning shows strong potential for universal image generation. We summarize four key findings: (1) it supports various in-domain tasks with reduced task ambiguity (Fig. 1); (2) it generalizes to unseen tasks (Fig. 2, Fig. 8); (3) as an unseen strategy for task unification, it can integrate multiple sub-tasks into a single step and generate intermediate results (Fig. 3); (4) it enables reverse generation, i.e., inferring a set of conditions from a given target (Fig. 9). While prior works [xiao2024omnigen, bar2022visual, zhang2023what, balazevic2023towards, Wang_2023_CVPR, liu2023unifying, alayrac2022flamingo] have also explored in-context learning in vision, they are largely constrained to specific domains (such as dense prediction or style transfer [Wang_2023_ICCV, zhu2024unleashing]), or simplified generation settings involving only one condition and one target image [liu2023unifying, sun2023imagebrush].
@@
From the perspective of task distribution, visual tasks are inherently sparse compared to those in natural language processing because task-specific datasets [xiao2024omnigen, zhou2017scene] for different tasks have minimal overlap [Yun_2023_ICCV, Kokkinos_2017_CVPR, Ghiasi_2021_ICCV]. Such sparse task learning isolates the knowledge of each task and limits the model from learning shared features across tasks. Moreover, the weak correlations between tasks hinder knowledge transfer and adaptability to new tasks. However, existing works in multi-task learning [ruder2017overview, Kendall_2018_CVPR, fifty2021efficiently, chen2018gradnorm] have verified the benefits of overlapping knowledge across related tasks. To alleviate the sparsity of visual tasks, we introduce a graph-structured dataset, Graph200K, where each image is associated with annotations spanning five meta-tasks, i.e., conditional generation [zhang2023adding], IP preservation [ye2023ip-adapter], style transfer [Zhang_2023_inst], image editing [wei2024omniedit], and restoration [yu2024scaling]. By combining different conditions, we train the model with a variety of tasks that overlap with each other. Given this highly overlapping and compact task space, our dataset significantly increases task density, allowing the model to learn shared and transferable knowledge more effectively.
@@
For the architecture design, it is essential to 1) accommodate flexible task formats [lhhuang2024iclora, xiao2024omnigen, le2024diffusiongenerate], ensuring seamless in-context learning, and 2) remain compatible with state-of-the-art models [zhuo2024luminanext, flux2024] to fully leverage their strong generative priors. In this work, we find that the state-of-the-art image infilling model [flux2024] has a consistent objective with our in-context learning based universal generative formulation. Specifically, we concatenate all input and output images together, where the objective of a task is to fill the output area. This alignment enables us to build our model upon advanced general-purpose infilling models without additional modifications, achieving powerful universal generation capabilities with minimal data and training costs.
@@
In this work, we propose a universal image generation framework, VisualCloze, which fine-tunes FLUX.1-Fill-dev [flux2024] with interrelated tasks sampled from Graph200K to learn transferable knowledge and support visual in-context learning. As the number of in-context examples increases, we observe enhanced performances and reduced task confusion, enabling the model to support a broad spectrum of in-domain tasks, including conditional generation, image restoration, editing, style transfer, IP-preservation, and their combinations. On unseen tasks, the model also shows a certain degree of generalization ability, as shown in Fig. 2. In summary, our main contributions are as follows:
@@ 2.1 Image Generation
Recent advances in text-to-image generation have achieved remarkable performance, largely driven by the development of autoregressive models [sun2024autoregressive, yu2022scaling, liu2024lumina-mgpt] and diffusion models [ddpm, gao2024mdtv2maskeddiffusiontransformer, Peebles_2023_ICCV, dhariwal2021diffusion, Rombach_2022_CVPR, esser2024scalingrectifiedflowtransformers, albergo2023building, lipman2023flow, liu2022flow]. Among these, rectified flow transformers [flux2024, esser2024scalingrectifiedflowtransformers, gao2024lumina, zhuo2024luminanext] have shown great training efficiency and overall performance. Building on these foundational models, diverse applications have emerged, such as conditional generation [zhang2023adding], style transfer [wang2024instantstyle], and personalized generation [li2023photomaker]. More recently, universal models that address various tasks [zhao2025diceptiongeneralistdiffusionmodel, mao2025ace++, le2024diffusiongenerate] have been explored. For example, unified models like OmniGen [xiao2024omnigen] leverage large vision language models to consolidate multiple tasks into a single framework. Similarly, UniReal [chen2024UniReal] unifies image generation tasks as discontinuous video generation. However, they still face issues such as over-reliance on language instructions, isolation and sparsity of visual tasks, and architecture design accommodating flexible task formats. To address these issues, we propose a universal image generation framework that unifies generation tasks as image infilling. Through visual in-context learning and our Graph200K dataset that constructs a denser task space to learn transferable knowledge, our method alleviates ambiguity to support a diverse set of in-domain tasks and generalizes to tasks unseen during training.
@@ 2.2 Visual In-context Learning
Along with the emergence of large language models, such as GPT-3 [brown2020language], in-context learning [dong2024surveyincontextlearning] has been an effective approach to allow the language model to understand and perform complex tasks given a few demonstrations. Early works [imageanalogies, hertzmann2001algorithms] in vision modality propose image analogies to create an image filter from examples automatically. In recent years, leveraging inpainting model [bar2022visual, zhang2023what, balazevic2023towards], masked image modeling [Wang_2023_ICCV, Wang_2023_CVPR, liu2023unifying], or vision-language model [alayrac2022flamingo, zhou2024visualincontextlearninglarge], visual in-context learning is proposed to handle more tasks. However, they mainly focus on dense prediction [zhu2024unleashing, Sheng_2024_CVPR, sun2023exploringeffectivefactorsimproving] or visual understanding [wang2024visincontext]. OmniGen [xiao2024omnigen] also leverages in-context learning to generalize to unseen domains, e.g., segmenting unseen concepts when the model has learned the segmentation task during training. However, it mainly focuses on simple tasks of dense prediction, and the gap between the unseen and training domains is still limited. Some recent works [liu2023unifying, sun2023imagebrush, wang2023incontext, lai2024unleashing] extend visual in-context learning to image generation, but they are still limited by simple tasks such as conditional generation and dense prediction. Moreover, the sparsity of visual tasks makes it difficult for models to learn transferable and overlapping knowledge across tasks, limiting the generation ability of in-context learning. In contrast, we introduce a graph-structured dataset that supports interrelated tasks and thus constructs a more dense task space, promoting the model to learn shared and transferable knowledge and enhance its adaptability.
@@ 3 Dataset
Recent works [mao2025ace++, lhhuang2024groupdiffusion, xiao2024omnigen] have made great progress in unified image generation. However, their generalization to unseen tasks remains highly limited. We partially attribute this issue to the sparsity and isolation of visual tasks, hindering the model from learning shared features across tasks and handling unseen ones. Moreover, weak correlations between tasks further hinder knowledge transfer, restricting the adaptability of models. Therefore, increasing task density or strengthening task inter-relations helps improve the generalization ability of models via a compact task distribution. In this paper, we take the Subject200K [tan2024ominicontrol] dataset as a starting point and construct our Graph200K dataset by augmenting each image with 49 types of annotations spanning five meta-tasks. This enriched annotation space enables flexible construction of a wide range of related tasks by sampling and combining arbitrary subsets of annotations across different meta-tasks, as illustrated in Fig. 4.
@@ 3.2 Dataset Construction
For convenience, we inherit subject-driven data from the Subjects200K [tan2024ominicontrol]. Additionally, 32 different degradations are applied online to the images to acquire restoration data. We summarize the data construction methods in this section for the remaining three tasks.
Conditional generation. Each image is paired with 12 distinct conditions generated by specialized models, including canny edges [canny1986computational], HED edges [xie15hed], Hough lines [gu2021realtime], semantic segmentation maps [uniformer], depth maps [depth_anything_v2], shape normal maps [xu2023unifying], and human keypoints [openpose], following ControlNet [zhang2023adding]. This work extends the conditions by incorporating SAM2 [ravi2024sam2] masks, foreground segmentation, and open-world boxes and masks. The foreground segmentation, derived from the RMBG [BiRefNet], supports diverse tasks such as inpainting and foreground extraction. Open-world bounding boxes are generated through the grounding caption capability of Qwen2-VL [Qwen2VL], which are processed using SAM2 [ravi2024sam2] to produce corresponding masks.
Style transfer. We transfer the style of images according to reference in both semantic-variant and semantic-invariant settings. Specifically, the semantic-invariant transfer adopts InstantStyle [wang2024instantstyle] to preserve the semantic content, while the semantic-variant transfer relies on FLUX.1-Redux-dev [flux2024], using the style embeddings and depth as conditions. For each image, we randomly generate five stylized versions. Mixing the two tasks pushes the model to follow the in-context examples better to avoid ambiguity.
Image editing. We design two types of editing tasks, including background-variant and background-invariant editing. The background-invariant editing begins with localizing the subjects. Then, we leverage a large vision-language model, Qwen2-VL [Qwen2VL], to modify the image caption with a new object that replaces the original subject. The image, with the subject masked, is subsequently processed by the FLUX.1-Fill-dev [flux2024] inpainting model to integrate the alternative object into the masked region. The above operation is repeated five times to enrich the dataset. For background-variant editing, the difference lies in the last step, which utilizes FLUX.1-Redux-dev [flux2024] with depth as the condition and the modified caption as the text prompt.
@@ 3.3 Other Data
To further expand the range of tasks and enhance the generalization ability of models, we incorporate several open-source datasets during training, including VITON-HD [choi2021viton] for virtual try-on and PhotoDoodle [huang2025photodoodlelearningartisticimage] for artistic image editing. For image editing tasks, we also extend the dataset with OmniEdit [wei2024omniedit]. Specifically, two sub-tasks, i.e., object addition and removal, are used for training. The other editing tasks, such as attribute modification and environment change, are treated as unseen tasks to assess the generalization ability of the trained model. Furthermore, we leverage a portion of high-quality internal data, covering tasks of the drawing process [paintsundo] and multi-view generation [huang2024mvadapter].
@@ 4.1 Visual In-context Learning
Language instructions are usually used to specify the generation definition to handle multiple visual generation tasks with a single generative model. However, due to the gap between vision and language, the text comprehension ability of image generation models remains limited. This issue leads to task confusion [lin2024pixwizard] in existing universal generative models and weak generalization to unseen tasks. Inspired by the success of few-shot learning on large language models [brown2020language], we recognize that visual context may serve as a more friendly task instruction for visual generative models, given their superior visual understanding capabilities.
Therefore, in this paper, we re-propose visual in-context learning to build a universal and generalizable image generation system. For the sake of description, here we assume the image input-output of arbitrary conditional generation task as a query consisting of L1L-1 condition images and a blank target $\varnothing$ to be completed by the model, i.e., ${X}={\rm concat}(\{x_{1},\dots,x_{L-1},\varnothing\})$. In Sec. 5.1, we demonstrate that our method can be extended to more general scenarios, where it can generate images at arbitrary positions and in any quantity rather than just the single image at the end of the query. During training, we randomly provide up to $C$ in-context examples, each containing $L$ images as the query. This strategy ensures the generalization ability of models across different numbers of in-context examples. In our experiments, we show that providing in-context examples as task demonstrations not only helps alleviate task confusion and boost model performance across in-domain tasks [lin2024pixwizard], but also enhances the generalization ability on unseen tasks.
@@ 4.2 Unified Multi-task Framework
Unlike previous visual in-context learning methods that primarily focus on scenarios with a single image condition and a single context [liu2023unifying, sun2023imagebrush], in this work, we aim to construct a unified framework capable of handling varying numbers of conditions and contexts, allowing for flexible adaptation to diverse tasks. For ease of description, we first assume all images processed by the model share the same size, $W\times H$, and we extend to the scenario with mismatched aspect ratios at the end of this section. In this way, given $C$ in-context examples and the query, each containing $L$ images, all images can be concatenated into a complete grid-layout image with a size of $(L\times W,(C+1)\times H)$. Then, the model can complete a task by infilling the target grids based on the surrounding context, akin to solving visual cloze puzzles. Therefore, we build our unified framework, VisualCloze, based on the general image infilling architecture capable of handling multiple resolutions.
@@
where $M(i,j)=1$ indicates that the pixel will be masked and generated by the infilling model. Equ. (2) masks the region in the last row and column, i.e., the target image. During training, we also randomly mask one of the first L1L-1 grids with a probability of 0.5, promoting reverse generation shown in Sec. 5.1. For the inference stage, we can crop $\hat{X}$ to obtain the target image easily.
Aligned optimization objective. A key benefit of this design is that our VisualCloze formulation shares a highly consistent objective with general image infilling models without architectural modifications or explicit input conditions. This consistency allows us to directly fine-tune advanced image infilling models using the newly constructed dataset while maximizing the utilization of the prior knowledge of foundation models. In contrast, existing task-specific models often require introducing additional learnable modules [li2023photomaker, wei2024omniedit] or adapting to extra condition inputs [tan2024ominicontrol], which may compromise the native capabilities of the model.
@@
Positional embedding. In the preceding section, all images are concatenated into a grid-layout image and we can apply positional embedding (i.e., RoPE [ropepaper]) on this large image. However, a potential limitation lies in composing a grid image from in-context examples with varying aspect ratios. To overcome this issue, we leverage the 3D-RoPE in Flux.1-Fill-dev to concatenate the query and in-context examples along the temporal dimension, as shown in Fig. 5, effectively overcoming this issue without introducing any noticeable performance degradation.